# Memory in the Burst Occurrence of Repeating FRBs

### v1

@@ 2 Results
Table 1: Values of the parameter $\gamma$ of the waiting time distribution $f(\lambda|\lambda_{0})$, preconditioned or not. $q_{1}$ and $q_{4}$ means $\lambda_{0}$ is from the first and the last part of sorted data, while $q_{0}$ is the unconditional case.
FRB 121102 FRB 20220912A FRB 20201124A(A) FRB 20201124A(B) Average
$q_{1}$ 2.40 3.07 2.55 3.88 $2.98\pm 0.65$
$q_{4}$ 2.08 2.06 2.13 2.07 $2.09\pm 0.41$
$q_{0}$ 2.22 2.63 2.40 3.04 $2.57\pm 0.38$
To quantify the impact of different preceding condition $\lambda_{0}$ on subsequent burst events, we proceed to calculate the average value of the following burst waiting time distribution, denoted as $\hat{\lambda}(\lambda_{0})$. From Eq. (1), we obtain a fluence independent functional form $\hat{\lambda}(\lambda_{0})\sim\bar{\lambda}g(\lambda_{0}/\bar{\lambda})$. Fig. 3 shows the ratio $\hat{\lambda}(\lambda_{0})/\bar{\lambda}$ as a function of $\lambda_{0}/\bar{\lambda}$. Due to the strong memory effect in the system, the value of $\hat{\lambda}/\bar{\lambda}$ is significantly below one when $\hat{\lambda}/\bar{\lambda}$ is much below one, and notably above one when $\hat{\lambda}/\bar{\lambda}$ is much above one. This implies that shorter waiting times are more likely to be followed by shorter ones, and longer waiting times by longer ones. For comparison, when the same preceding condition $\lambda_{0}$ is applied after random shuffling, the value of $\hat{\lambda}/\bar{\lambda}$ is approximately 1, indicating that the memory effect has been disrupted.
@@
The obtained $\hat{\lambda}(\lambda_{0})$ reflects the average time to wait for the next FRB burst under different preceding condition values $\lambda_{0}$. If we want to know more detailed temporal information about the next burst when the time interval between the previous two bursts is $\lambda_{0}$, this can be obtained by calculating the probability of another burst occurring within a certain time (or risk function), denoted as $R(t|\lambda_{0})$. Fig. 4 compares the probability distribution of the timing of the next burst when the time interval $\lambda_{0}$ between the previous two bursts is either minimal ($\in q_{1}$) or maxmal ($\in q_{4}$). For FRB 121102 (Fig. 4(a)) and FRB 20220912A (Fig. 4(b)), the value of $\lambda_{0}$ has a significant impact on the timing of the next burst within a certain time range. As a comparison, for the same burst source FRB 20201124A(A) and FRB 20201124A(B) during different active periods, $R(t|\lambda_{0})$ is relatively less affected by the magnitude of the preceding condition values $\lambda_{0}$, but maintains the property that a smaller $\lambda_{0}$ is followed by a shorter time until the next FRB burst, and a larger $\lambda_{0}$ is followed by a relatively longer waiting time for the next burst. These characteristics are important in understanding the bursting nature of FRBs.
@@
Figure 4: The conditional probability of waiting time $R(t|\lambda_{0})$, exceeding a specified fluence threshold and separated by time $\lambda_{0}$, are succeeded by the next FRB within time t. The $\lambda_{0}$ comes from either $q_{1}$ or $q_{4}$ and thresholds are 0.1, 0.6 and 0.5 Jy ms for FRB 121102, FRB 20220912A and FRB 20201124A(B) and 1.0 Jy s for FRB 20201124A(A).
@@
From the preceding analysis, it is evident that the time intervals between the first two bursts in repeating FRBs exert a significant influence on the occurrence of subsequent bursts. A realistic question arises: if a time interval $t$ has elapsed since the last burst without a subsequent occurrence, what is the average waiting time until the next burst? To answer this, one must calculate the expected residual time $\hat{\lambda}(t|\lambda_{0})$. For the special case when $t=0$, $\hat{\lambda}(t=0|\lambda_{0})=\hat{\lambda}(\lambda_{0})$. Based on the definition of residual time and the known waiting time distribution (Eq. 2), one can deduce that
@@
where the integral is divergence when $\gamma$ is equal to $1.0$ and $2.0$.
@@
When there is no memory effect, $\hat{\lambda}(t|\lambda_{0})$ is independent of both $\lambda_{0}$ and $t$. Conversely, if a memory effect exists, a clear dependency will be observed. To quantitatively characterize the dependency of $\hat{\lambda}$ on $\lambda_{0}$ and $t$, the following scenarios were considered: firstly, intervals $q_{+}$ and $q_{-}$ that are respectively greater and smaller than its median, and $\lambda_{0}$ is taken from intervals $q_{+}$ and $q_{-}$, denoted as $\lambda^{+}_{0}$ and $\lambda^{-}_{0}$. Additionally, $\hat{\lambda}(t)$ without any preceding condition is also considered; secondly, For comparison, $t$ is set to either 00 or $\bar{\lambda}/2$. The results are shown in Fig. 5, where sample A, B, C, D represent the datasets for FRB 121102, FRB 20220912A, FRB 20201124A(A), and FRB 20201124A(B), respectively. For the case of $t=0$ (Fig. 5(a)), since $\hat{\lambda}(0)=(\hat{\lambda}(0|\lambda^{-}_{0})+\hat{\lambda}(0|\lambda^{+}_{0}))/2$, the results show that $\hat{\lambda}(0|\lambda^{-}_{0})$ (Blue square) and $\hat{\lambda}(0|\lambda^{+}_{0})$ (Blue left triangle) are approximately symmetrical about $\hat{\lambda}(0)$ (Green solid circle) $=1$, and due to the memory effect, $\hat{\lambda}(0|\lambda^{+}_{0})$ is significantly greater than $\hat{\lambda}(0|\lambda^{-}_{0})$. Similar memory effects as shown in Fig.5(a) are also evident in Fig. 5(b), i.e., $\hat{\lambda}(\frac{\bar{\lambda}}{2}|\lambda^{+}_{0})$ is significantly greater than $\hat{\lambda}(\frac{\bar{\lambda}}{2}|\lambda^{-}_{0})$. Furthermore, due to the memory effect, as the time elapsed is $t=\lambda/2$ since the last burst was introduced, the symmetry between the residual times as shown in Fig. 5(a) is broken due to the arising of $\bar{\lambda}$, which is mostly contributed from $q_{+}$.
@@
Figure 5: The conditional residual time $\hat{\lambda}(t|\lambda_{0})$. The $\lambda_{0}$ is either from $q_{+}$ (above median of entire data) or $q_{-}$ (below median of entire data). Two cases are considered for (a) $t=0$ and (b) $t=\bar{\lambda}/2$. Datasets A, B, C, and D correspond to FRB 121102, FRB 20220912A, FRB 20201124A(A), and FRB 20201124A(B), respectively. Fluences thresholds are the same as in Fig. 4.
@@
As illustrated in Fig. 5, after a time interval $t$ has elapsed since the last burst, the expected residual time to the next burst is generally greater than when $t=0$. This can be directly inferred from Eq. 4. Generally, as $t$ increases, $\hat{\lambda}(t|\lambda_{0})$ also monotonically increases, with the rate of increase influenced by the value of $\gamma$. This behavior contrasts with the memoryless nature of the Poisson process (Sornette & Knopoff, 1997; Corral, 2005; Livina et al., 2005).
@@ 3 Discussion
Contrary to intuitive view, the probability of next burst occurring does not increase with the time $t$ elapsed since the last burst; rather, it decreases and the expected residual time to the next burst $\hat{\lambda}$ is also increased. This serves as a more direct characterization of temporal clustering, where waiting times exhibit a tendency to cluster together.
@@
An intuitive explanation for this phenomenon shown that when a significant amount of time has elapsed since the last FRB burst, the system enters a enhanced ’period of inactivity,’ during which the likelihood of an extended waiting time increases. It is important to note, however, that these periods of inactivity are not fundamentally distinct from other waiting times, as they are all governed by the same continuous distribution. Furthermore, the fluence of the subsequent burst does not increase with the waiting time but appears to be independent of it.
@@
Recent research, as referenced in (Aschwanden & McTiernan, 2010), offers a generalized understanding of nonstationary Poisson processes, which maybe helpful in interpreting the results. Specifically, when the variability of the burst rate follows a gradual pattern and can be described by piecewise exponential functions, the exponent $\gamma$ in the waiting time distribution approximates 3.0. Conversely, if the burst rate variability exhibits sharp, intermittent spikes like to $\delta$-functions, the exponent $\gamma$ in the waiting time distribution tends toward 2.0. This behavior, commonly referred to as ”clusterization”, finds also in earthquake and solar flare statistics (Corral, 2004, 2005; Livina et al., 2005; Baiesi et al., 2006; Aschwanden & McTiernan, 2010), which serves as indirect evidence for non-trivial space-time correlations among events. Consequently, the value of the exponent $\gamma$ serves as a critical indicator for determining whether the burst rate undergoes gradual or intermittent fluctuations. As a result, the bursts themselves are not periodic, the activity may occur within periodic windows (Wei et al., 2022). Some recent studies have sought to distinguish between periodic windowed behavior and mere event clusters through time clustering analysis (Denissenya et al., 2021).
@@
In this study, the analysis of four datasets from three FRB sources reveals that the value of $\gamma$ in the waiting time distribution (Eq. (2)) ranges approximately between $2.2$ and $3.0$ (Table 1). For cases where $\gamma$ is close to $2.0$, we made an estimate based on Eq. (4), when $\beta=1.0,\gamma=2.1$, the expected residual time to the next burst $\hat{\lambda}$ is roughly 10 times the magnitude of the elapsed time $t$ after scaling. Here, the actual value of each scale $t$ is related to the average waiting time under different fluence thresholds. If $\beta$ remains constant and $\gamma$ decreases to $2.01$, then $\hat{\lambda}$ will be roughly $100$ times $t$. As $\gamma$ gradually approaches $2.0$, $\hat{\lambda}$ will become larger. The three burst sources analyzed in this paper have very different burst environments and properties. However, the statistical analysis results of these three repeating FRBs with the largest sample sizes indicate that the scale invariant waiting time distribution and the clustering and memory of bursts are universal. This may be well used to explain the burst event characteristics of repeating FRBs that have been active in the past, as well as to explain the burst event characteristics that have only been observed once but may be repeating bursts. Although the existence of true one-offs is not ruled out, Eq. (4) has also shown that when $\gamma=2.0$, the expected residual time diverges and tends to be infinite, and this corresponds to the case of spikes like $\delta$ function of time dependent burst rate (Aschwanden & McTiernan, 2010), which mimics a true one-off burst from an observational perspective. Therefore Eq. (2) actually provides a unified description of the statistical behavior of the waiting time distribution, satisfying this equation for both repeating FRBs and one-off FRBs.
@@
The complexity of waiting time between FRBs suggests that the physical models proposed thus far may not fully capture the complexity of the phenomenon. Currently, over 700 FRBs have been observed, of which more than 60 exhibit repeating bursts (Xu et al., 2023). Some analyses suggest that one-off and repeating fast radio bursts have some statistical differences (Chen et al., 2022; Cui et al., 2020). There are also studies discussing whether repeating and non-repeating FRBs have the same burst mechanisms (Caleb et al., 2019; Zhang, 2020, 2023), and partially explaining the statistical differences between one-off and repeating FRBs from an observational perspective (Iwazaki, 2021). However, more research is needed to determine the precise mechanisms that lead to the intermittent nature of FRBs and the extended periods between bursts. The ubiquity of this behavior implies the presence of a straightforward mechanism. In this mechanism, as time progresses, the variable responsible for triggering the burst increasingly deviates from the burst threshold, on average. These ’fluctuations’ in the variable effectively preserve the memory of the last event within the system for extended periods, thereby contributing to the observed absence of time dependent aging. It is important to emphasize that these findings are in fact model independent, meaning they do not rely on any preconceived model of burst occurrence. Bak et al. (Bak, 1996; Bak et al., 2004) suggests that the observed scaling law in burst events reflects a complex spatiotemporal pattern. This may be interpreted as intermittent energy release from a self-organized system , which is consistent with the theoretical constructs of critical phenomena. Analogous memory effects, attributable to long-term correlations, have also been recently identified in the statistical analysis of two FRBs (Wang et al., 2023), which address the self-similarity inherent in these processes, the framework of SOC currently offers the most compelling approach.

### v2

@@ 2 Method and Results
Table 1: Best-fitting values of the parameter $\gamma$, its $1\sigma$ errors and reduced Chi-square of the waiting time distribution $f(\lambda|\lambda_{0})$, preconditioned or not. $q_{1}$, $q_{2}$, $q_{3}$ and $q_{4}$ means $\lambda_{0}$ is from the first, second, third and the last part of sorted data, while $q_{0}$ is the unconditional case.
FRB 20121102A FRB 20220912A FRB 20201124A(A) FRB 20201124A(B)
$\gamma$           $\chi^{2}$ $\gamma$           $\chi^{2}$ $\gamma$           $\chi^{2}$ $\gamma$           $\chi^{2}$ Mean of $\gamma$
$q_{1}$ $4.80^{+0.37}_{-0.52}$   1.38 $4.33^{+2.38}_{-1.18}$   0.77 $3.95^{+0.97}_{-0.57}$   2.31 $4.34^{+0.58}_{-0.71}$   0.48 $4.26$
$q_{2}$ $4.38^{+0.23}_{-0.19}$   1.23 $4.27^{+1.80}_{-0.98}$   0.60 $3.06^{+0.57}_{-0.37}$   1.12 $3.19^{+0.62}_{-0.41}$   0.49 $3.73$
$q_{3}$ $3.24^{+0.31}_{-0.25}$   1.05 $2.66^{+0.46}_{-0.36}$   0.78 $2.92^{+0.37}_{-0.27}$   0.93 $2.76^{+0.17}_{-0.27}$   0.52 $2.90$
$q_{4}$ $2.76^{+0.51}_{-0.24}$   0.75 $2.61^{+0.69}_{-0.15}$   0.65 $2.34^{+0.20}_{-0.15}$   0.59 $2.08^{+0.31}_{-0.23}$   0.80 $2.45$
$q_{0}$ $3.72^{+0.32}_{-0.23}$   2.17 $3.34^{+0.29}_{-0.24}$   0.33 $2.78^{+0.12}_{-0.11}$   2.39 $2.65^{+0.21}_{-0.17}$   0.67 $3.12$
[…]
@@
The obtained $\hat{\lambda}(\lambda_{0})$ reflects the average time to wait for the next FRB burst under different preceding condition values $\lambda_{0}$. If we want to know more detailed temporal information about the next burst when the time interval between the previous two bursts is $\lambda_{0}$, this can be obtained by calculating the probability of the next burst occurring within a certain time $t$ or the cumulative distribution of $P(\lambda)$ (Eq. 1), denoted $R(t|\lambda_{0})$. Here, $t$ is a specific moment and the integral is conducted in the range of $(0,t)$. This normalized probability reflects the distribution probabilities of the subsequent burst intervals in a different time range ($<t/\bar{\lambda}$) when the precondition is [$q_{1},q_{2},q_{3},q_{4}$]. As $t$ becomes sufficiently large, the probability tends towards 1. Fig. 4 compares the probability distribution of the timing of the next burst when the time interval $\lambda_{0}$ is between the previous two bursts of $[q_{1},q_{2},q_{3},q_{4}]$. For FRB 20121102A (Fig. 4(a)) and FRB 20220912A (Fig. 4(c)), the value of $\lambda_{0}$ clearly influences the timing of the next burst. This impact is particularly evident in the results $q_{1}$ and $q_{4}$, where the differences are especially pronounced. For FRB 20201124A(A) (Fig. 4(b)) and FRB 20201124A(B) (Fig. 4(d)) during different active periods, $R(t|\lambda_{0})$ is relatively less affected by the magnitude of the values of the preceding conditions $\lambda_{0}$.
We conducted a two-sample K-S test comparing the differential distributions $P(\lambda|\lambda_{0})$ where the preceding condition values $\lambda_{0}$ come from $q_{1}$ and $q_{4}$. The K-S statistics and P-values for FRB 20121102A, FRB 20220912A, FRB 20201124A(A) and FRB 20201124A(B) are as follows: (0.38, 0.14), (0.39, 0.11), (0.22, 0.78) and (0.37, 0.19). All p-values are greater than the 0.05 threshold, indicating that there are no statistically significant differences. However, the small size and discrete nature of the binned data likely contribute to the increase in P-values, potentially obscuring subtle differences between distributions that the K-S test might miss. Therefore, a larger sample is potentially needed to further investigate and perhaps arrive at a significant p-value. For the results of FRB 20201124A(A), it could be associated with the instability of these distributions when dealing with limited statistics, with the expectation of achieving improved statistical results as more observational data become available. Naturally, this quantity $R(t|\lambda_{0})$ represents only the average effect of the distribution function in specific contexts and, to some extent, may also highlight the complexity of the FRB waiting time distribution, which may be crucial to understanding the bursting behavior of FRBs.
@@
Figure 4: The conditional probability of waiting time $R(t|\lambda_{0})$, exceeding a specified fluence threshold for four FRBs datasets respectively, and separated by previous time interval $\lambda_{0}$, are succeeded by the next FRB within time t. The $\lambda_{0}$ comes from either $q_{1}$, $q_{2}$, $q_{3}$ or $q_{4}$ and thresholds are 0.1, 0.6, 0.5 Jy ms and 1.0 Jy s for FRB 20121102A, FRB 20220912A, FRB 20201124A(B) and FRB 20201124A(A).
@@
From the preceding analysis, it is statistically evident that the time intervals between the first two bursts in repeating FRBs generally exert an evident influence on the occurrence of subsequent bursts. A realistic question arises: If a time $t$ has elapsed since the last burst without a subsequent occurrence, what is the average waiting time until the next burst? To answer this, one must calculate the expected residual time $\hat{\lambda}(t|\lambda_{0})$. For the special case $t=0$, which means that there is no time elapsed since the last burst, $\hat{\lambda}(t=0|\lambda_{0})=\hat{\lambda}(\lambda_{0})$. Based on the definition of residual time and the known waiting time distribution (Eq. 2), one can deduce that
@@
where the integral is conducted under the condition that $\gamma>2.0$.
@@
When there is no memory effect, $\hat{\lambda}(t|\lambda_{0})$ is independent of both $\lambda_{0}$ and $t$. Conversely, if a memory effect exists, a clear dependency will be observed. The uncertainty in $\hat{\lambda}(t|\lambda_{0})$ arises from the statistical error in the number of waiting time data (Lin & Sang, 2019; Wei et al., 2021) measured at $t=0$ or other values. This uncertainty is then propagated into the calculation of the residual time. To quantitatively characterize the dependency of $\hat{\lambda}$ on $\lambda_{0}$ and $t$, the following scenarios were considered: First, the entire waiting time dataset is divided into $q_{+}$ and $q_{-}$, respectively, which is greater and smaller than its median, and $\lambda_{0}$ is taken from the intervals $q_{+}$ and $q_{-}$, denoted as $\lambda^{+}_{0}$ and $\lambda^{-}_{0}$. Furthermore, $\hat{\lambda}(t)$ without any preceding condition is also considered; Second, for comparison, $t$ is set to 00 or $\bar{\lambda}/2$. The results are shown in Fig. 5, where the samples are from the datasets of FRB 20121102A, FRB 20220912A, FRB 20201124A(A) and FRB 20201124A(B), respectively. For the case of $t=0$ (Fig. 5(a)), since $\hat{\lambda}(0)=(\hat{\lambda}(0|\lambda^{-}_{0})+\hat{\lambda}(0|\lambda^{+}_{0}))/2$, the results show that $\hat{\lambda}(0|\lambda^{-}_{0})$ (solid green triangle) and $\hat{\lambda}(0|\lambda^{+}_{0})$ (solid red triangle) are approximately symmetrical to $\hat{\lambda}(0)$ (solid blue circle) $=1$ (relative to $\bar{\lambda}$), and due to the memory effect, $\hat{\lambda}(0|\lambda^{+}_{0})$ is significantly greater than $\hat{\lambda}(0|\lambda^{-}_{0})$. The statistical significance of these differences in this case is $(4.5,2.9,4.1,4.2)$ for FRB 20121102A, FRB 20220912A, FRB 20201124A(A) and FRB 20201124A(B), respectively. Similar memory effects as shown in Fig. 5(a) are also evident in Fig. 5(b), that is, $\hat{\lambda}(\frac{\bar{\lambda}}{2}|\lambda^{+}_{0})$ is generally greater than $\hat{\lambda}(\frac{\bar{\lambda}}{2}|\lambda^{-}_{0})$. In this case, the significance of these differences is (0.8, 1.5, 2.9, 3.2). Compared to the situation at $t=0$, the reduction in significance may primarily be due to the increased statistical fluctuations caused by the reduced sample size. Furthermore, due to the memory effect, since the time elapsed is $t=\bar{\lambda}/2$ for the last burst was introduced, the symmetry between the residual times as shown in Fig. 5(a) is broken due to the arising of $\bar{\lambda}$, which is mostly contributed from $q_{+}$.
@@
Figure 5: The conditional residual time $\hat{\lambda}(t|\lambda_{0})$. The $\lambda_{0}$ is either from $q_{+}$ (above median of entire data), $q_{-}$ (below median of entire data) or $q_{0}$ (the unconditional case). Two cases of $\hat{\lambda}(t|\lambda_{0})$ for elapsed time t are considered for (a) $t=0$ and (b) $t=\bar{\lambda}/2$. In the two cases all FRBs studied exhibit $\hat{\lambda}(t|\lambda_{0}^{+})>\hat{\lambda}(t|\lambda_{0}^{-})$, and are symmetrically distributed around $\hat{\lambda}(t)$ at $t=0$. Additionally, relative to t = 0, the residual time $\hat{\lambda}(t|\lambda_{0})$ at $t=\frac{\bar{\lambda}}{2}$ generally increases, regardless of whether $\lambda_{0}$ comes from $q_{+}$, $q_{-}$ or $q_{0}$. Fluences thresholds are the same as in Fig. 4.
@@
As illustrated in Fig. 5, after a time interval $t$ has elapsed since the last burst, the expected residual time to the next burst is generally greater than when $t=0$. This can be directly inferred from Eq. 4. Generally, as $t$ increases, $\hat{\lambda}(t|\lambda_{0})$ also increases monotonically, with the rate of increase influenced by the value of $\gamma$. Especially, for the case of $\gamma$ close to $2.0$ the $\hat{\lambda}(t|\lambda_{0})$ is divergent. This behavior contrasts with the memoryless nature of the Poisson process (Sornette & Knopoff, 1997; Corral, 2005; Livina et al., 2005).
@@ 3 Discussion
Contrary to intuitive view, the probability that the next burst will occur does not increase with the time $t$ elapsed since the last burst; rather, it decreases, and the expected residual time to the next burst $\hat{\lambda}$ also increases. This serves as a more direct characterization of temporal clustering, where waiting times exhibit a tendency to cluster together.
@@
An intuitive explanation for this phenomenon has shown that when a significant amount of time has elapsed since the last FRB burst, the system enters an enhanced “period of inactivity”, during which the likelihood of an extended waiting time increases. However, it is important to note that these periods of inactivity are not fundamentally distinct from other waiting times, as they are all governed by the same continuous distribution. Furthermore, the fluence of the subsequent burst does not increase with the waiting time but appears to be independent of it.
@@
Recent research, as referenced in Aschwanden & McTiernan (2010), offers a generalized understanding of non-stationary Poisson processes, which may be helpful in interpreting the results. Specifically, when the variability of the burst rate follows a gradual pattern and can be described by piecewise exponential functions, the exponent $\gamma$ in the waiting time distribution approximates 3.0. Conversely, if the burst rate variability exhibits sharp, intermittent spikes akin to $\delta$ functions, the exponent $\gamma$ in the waiting time distribution tends toward 2.0. This behavior, commonly referred to as “clusterization”, is also found in earthquake and solar flare statistics (Corral, 2004, 2005; Baiesi et al., 2006; Aschwanden & McTiernan, 2010), which serves as indirect evidence for non-trivial space-time correlations among events. Consequently, the value of the exponent $\gamma$ serves as a critical indicator to determine whether the burst rate undergoes gradual or intermittent fluctuations. As a result, the bursts themselves are not periodic; the activity may occur within periodic windows (Wei et al., 2022). Some recent studies have sought to distinguish between periodic windowed behavior and mere event clusters through time cluster analysis (Denissenya et al., 2021). More in-depth studies have shown that a similar memory effect is also observed in earthquakes (Livina et al., 2005). If repeating FRBs exhibit burst characteristics that are similar, or partially similar, to those observed in earthquakes or solar flares, the phenomena of “clusterization” and “memory effects” observed in these events could be more easily comprehensible. Indeed, current research (Du et al., 2024) indicates that the scaling properties of repeating FRBs over longer time intervals (specifically, those exceeding approximately 1 s) are similar to those observed in solar flares, which could indicate a shared nature in their fundamental processes. In contrast, FRBs characterized by shorter intervals, that is, less than about 1 second, exhibit characteristics that are more similar to those associated with the dynamics of earthquakes (Totani & Tsuzuki, 2023).
@@
In this study, the analysis of four datasets from three FRB sources reveals the value of $\gamma$ in the waiting time distribution (Eq. (2)) ranges approximately between $2.65$ and $3.72$ (for $q_{0}$ in Table 1). For cases where $\gamma$ is close to $2.0$, we make an estimate based on Eq. (4), when $\beta=1.0,\gamma=2.1$, the expected residual time to the next burst $\hat{\lambda}$ is roughly 10 times the magnitude of the elapsed time $t$ after scaling. Here, the actual value of each scale $t$ is related to the average waiting time under different fluence thresholds. If $\beta$ remains constant and $\gamma$ decreases to $2.01$, then $\hat{\lambda}$ will be roughly $100$ times $t$. As $\gamma$ gradually approaches $2.0$, $\hat{\lambda}$ will become larger. The three burst sources analyzed in this paper have very different burst environments and properties. However, the statistical analysis results of these three repeating FRBs with the largest sample sizes indicate that the scale-invariant waiting time distribution and the clustering and memory of bursts are universal. This may be well used to explain the burst event characteristics of repeating FRBs that have been active in the past, as well as to explain the burst event characteristics that have only been observed once but may be repeating bursts. Although the existence of true one-offs is not ruled out, Eq. (4) has also shown that when $\gamma=2.0$, the expected residual time diverges and tends to be infinite, and this corresponds to the case of spikes like $\delta$ function of time dependent burst rate (Aschwanden & McTiernan, 2010), which mimics a true one-off burst from an observational perspective. Hence, Eq. (2) offers a comprehensive explanation of the statistical characteristics of the waiting time distribution, accommodating the behavior of both repeating and one-off FRBs.
@@
The complexity of waiting time between FRBs suggests that the physical models proposed thus far may not fully capture the complexity of the phenomenon. Currently, more than 700 FRBs have been observed, of which more than 60 exhibit repeating bursts (Xu et al., 2023). Some analyses suggest that one-off and repeating FRBs have some statistical differences (Chen et al., 2022; Cui et al., 2020). There are also studies discussing whether repeating and non-repeating FRBs have the same burst mechanisms (Caleb et al., 2019; Zhang, 2020, 2023; Kirsten et al., 2024), and partially explaining the statistical differences between one-off and repeating FRBs from an observational perspective (Iwazaki, 2021). However, more research is needed to determine the precise mechanisms that lead to the intermittent nature of FRBs and the extended periods between bursts. The ubiquity of this behavior implies the presence of a straightforward mechanism. In this mechanism, as time progresses, the variable responsible for triggering the burst increasingly deviates from the burst threshold on average. These “fluctuations” in the variable effectively preserve the memory of the last event within the system for extended periods, thereby contributing to the observed absence of time dependent aging. It is important to emphasize that these findings are in fact model independent, which means that they do not rely on any preconceived model of burst occurrence. Bak et al. (Bak, 1996; Bak et al., 2002) suggests that the observed scaling law in burst events reflects a complex spatiotemporal pattern. This may be interpreted as an intermittent energy release from a self-organized system , which is consistent with the theoretical constructs of critical phenomena. Analogous memory effects, attributable to long-term correlations, have also been recently identified in the statistical analysis of two FRBs (Wang et al., 2023), which address the self-similarity inherent in these processes, the framework of SOC currently offers the most compelling approach.
Our findings highlight general trends, but the limited sample size of independent repeating FRBs means that some exceptions may arise, and our conclusions should be interpreted with caution. The rarity and variability of FRBs underscore the need for more extensive datasets to validate and refine these preliminary insights. Advancements in detection technologies will likely provide the larger datasets needed to support our initial findings and enable more detailed studies of the characteristics and effects of FRB memory. Future observations will be crucial for verifying these trends, refining our understanding of FRB behavior, and uncovering significant patterns that may not yet be fully apparent. Continued research and data collection are essential for advancing our knowledge and confirming or challenging these early findings, ultimately deepening our understanding of these cosmic phenomena.